OpenStudy (anonymous):

Rewrite the expression without using grouping symbols. 7(x - 19) 7 + 7 • 19 7x - 26 7x - 19 7x - 7 • 19

OpenStudy (the_fizicx99):

D

OpenStudy (anonymous):

Thanks! I have another can you help?

OpenStudy (the_fizicx99):

Its 7*x and 7*-19 So 7x - 7*19

OpenStudy (anonymous):

Thanks so much

OpenStudy (the_fizicx99):

Yw

OpenStudy (anonymous):

Could help me with another one?

OpenStudy (the_fizicx99):

sure

OpenStudy (anonymous):

Rewrite the expression using grouping symbols. 29x - 13x x(13 -29) x(29 -13) 13(x - 29) 13(x + 29)

OpenStudy (the_fizicx99):

C x*29 - x*13

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

I have two more I need help with.

OpenStudy (the_fizicx99):

Kay

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Which shows how to use the distributive property to evaluate 6 • 78? 6 • (70 - 8) 6 • (70 + 2) 6 • (70 + 8) 6 • (70 - 2)

OpenStudy (the_fizicx99):

Which one do you think it is

OpenStudy (the_fizicx99):

Can't keep giving out answer lol

OpenStudy (anonymous):

I think its B

OpenStudy (the_fizicx99):

Distributive is when you multiply everything in the parenthesis: Ex: 5(x + 5) 5*x + 5*5 5x + 25

OpenStudy (anonymous):

Ok what next

OpenStudy (the_fizicx99):

But see that I used x and 5, so they weren't like terms. Ex: 5(10 + 5) 5(15) 5*15 = 75

OpenStudy (the_fizicx99):

Soo, look at your options 70 + 8 = ??

OpenStudy (anonymous):

70 + 8 = 78

OpenStudy (the_fizicx99):

You can even do its separately, 5 * 10 = 50 5*5 = 25 25 + 50 = 75

OpenStudy (the_fizicx99):

So, which one do you think it is now.

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

I may sound dumb, but I think it's D now.

OpenStudy (the_fizicx99):

xD not D, 70 - 2 = 68 6 * 68 isn't what they're asking you.

OpenStudy (anonymous):

I am kinda of a blonde, I don't understand math too well!

OpenStudy (anonymous):

6 • 68 = 408

OpenStudy (the_fizicx99):

xD no worries, I have patience. Soo 70 + 8 = 78 * 6 They're asking for 6 * 78

OpenStudy (anonymous):

Ohhhh ok

OpenStudy (anonymous):

6 • 78 = 468

OpenStudy (the_fizicx99):

Which option do you think its now. Btw they're not asking for the answer, just the set up

OpenStudy (anonymous):

It's between A & C

OpenStudy (the_fizicx99):

Why would you subtract 8 from 70, that would be 62 * 6 .. nope

OpenStudy (anonymous):

right...

OpenStudy (anonymous):

It's C

OpenStudy (the_fizicx99):

Yup, could you explain why you think its C? Cx

OpenStudy (anonymous):

Sorry my internet connection went out

OpenStudy (anonymous):

I think it's C because 70+8 = 78 and then that's where I end up at 6 • 78

OpenStudy (anonymous):

Am I right?

OpenStudy (the_fizicx99):

Yes